A painter can do a given job in 6 hours. Her helper can do the same job in 10 hours. The painter begins the work and after two hours is joined by the helper. In how many hours will they complete the job?
Answer by mangopeeler07(462)

A painter can do a given job in 6 hours. Her helper can do the same job in 10 hours. The painter begins the work and after two hours is joined by the helper. In how many hours will they complete the job?

After 2 hours, the painter has done 2/6 or 1/3 of the work, so how much time will it take to do the other 2/3 of the job?

1/6+1/10=x
x=16/60-------------------rate of work together

16/60y=2/3

y=2.5

It will take them 2.5 hours to finish the job.